7 part 3

Consider the following data on US GDP

7. Problems and Applications Q7 Consider the following data on U.S. GDP: Year 2013 1993 Nomlnal GDP (Blllions of dollars) 16,663 6,879 GDP Deflator (Base year 2009) 106.9 72.3 The growth rate of nominal GDP between 1993 and 2013 was

The growth rate of nominal GDP between 1993 and 2013 was ____ , and the growth rate of the GDP deflator between 1993 and 2013 was ____ .(Hint: The growth rate of a variable XX over an NN-year period is calculated as 100×((XfinalXinitial)(1N)−1)100×XfinalXinitial1N−1)

Measured in 2009 prices, real GDP was ___ billion in 1993 and ___ billion in 2013.

The growth rate of real GDP between 1993 and 2013 was ____ .

The growth rate of nominal GDP between 1993 and 2013 was____than the growth rate of real GDP.


media%2F16a%2F16a56048-bb5f-4ac9-a932-8d
media%2Ff76%2Ff76a5208-4e13-4af7-9ee7-10
media%2F1df%2F1dfbd165-b964-4c73-8589-2a
media%2F39b%2F39be3a68-1a77-4f06-a899-03
media%2Fff8%2Fff8d88f1-7ad2-49ff-9c1f-87
media%2F69a%2F69ae682e-8fff-462b-8248-5d

(Please make sure you use answers from the answer choices)
Answer #1

The growth rate of nominal GDP between 1993 and 2013 is 100×((Xfinal/Xinitial)^(1/N)?1)

= ((16663/6879)^(1/20) - 1 )*100

= (2.4^(1/20) - 1 )*100

= (1.045- 1 )*100

= 4.5%

the growth rate of the GDP deflator between 1993 and 2013 is 100×((Xfinal/Xinitial)^(1/N)?1)

= ((106.9/72.3)^(1/20) - 1 )*100

= (1.48^(1/20) - 1 )*100

= (1.02- 1 )*100

= 2%

Measured in 2009 prices, real GDP was ___ billion in 1993 and ___ billion in 2013. (2009 prices not given here)

real GDP

2013 = 1666300/106.9 = 15587.46

1993 = 687900/72.3 = 9514.52

The growth rate of real GDP between 1993 and 2013 is 100×((Xfinal/Xinitial)^(1/N)?1)

= ((15587.46/9514.52)^(1/20) - 1 )*100

= (1.64^(1/20) - 1 )*100

= (1.025- 1 )*100

= 2.5%

The growth rate of nominal GDP between 1993 and 2013 was (4.5%) higher than the growth rate of real GDP (2.5%).